OpenStudy (anonymous):

Whats the equation for a question as following: "What is the 20th term of the arithmetic sequence 4, –2, –8, "

OpenStudy (asnaseer):

This site might be helpful: http://www.mathsisfun.com/algebra/sequences-sums-arithmetic.html

OpenStudy (anonymous):

hmmmm @asnaseer could you help me with the question above then. I get that xn=a+d(n-1) but Im not sure how to plug stuff in:/

OpenStudy (asnaseer):

in the equation you just posted, we have:\[x_n=a+d(n-1)\]have your read that web page to understand what each of these symbols stand for?

OpenStudy (asnaseer):

for example, what does the "a" stand for?

OpenStudy (anonymous):

a=first term. So in this case "4."

OpenStudy (asnaseer):

correct - what about the other symbols? d = ? n = ?

OpenStudy (anonymous):

d=-6 n=?

OpenStudy (asnaseer):

do you know what "n" stands for?

OpenStudy (anonymous):

...its the term right? So 20?

OpenStudy (asnaseer):

perfect! :) yes, "n" stands for the n'th term - in your case you want the 20'th term so (as you said) n=20

OpenStudy (asnaseer):

so use the formula to wok out \(x_{20}\)

OpenStudy (anonymous):

S0 x20= 4+ -6(19)? x20=4-114 x20=-110?:)

OpenStudy (asnaseer):

perfect again! you have the right answer. :)

OpenStudy (anonymous):

Ha Thanks! Last thing, so I don't start a whole new questions: "What is the sum of a 38–term arithmetic sequence where the first term is 14 and the last term is 154?" whats d?:/

OpenStudy (asnaseer):

The code of conduct for this site states that you should post each question separately. This also helps others learn by them being able to see your question and delving in if they are interested.

OpenStudy (asnaseer):

I'll help you once you have posted it. :)